Solve the following recurrence relations for the initial conditions given. The population of Utopia increases 5 percent per year. In 2000 the population was 10,000. What was the population in 1970?



1
Expert's answer
2021-11-23T11:11:01-0500
"P(t)=P_0(1+r)^t"

Given "r=0.05, P(30)=10000"


"P_0=\dfrac{P(30)}{(1+r)^{30}}=\dfrac{10000}{(1+0.05)^{30}}=2314"

The population was 2314 in 1970.